OpenStudy (anonymous):

HELP Solve. 3x2 = -21x - 15

OpenStudy (midhun.madhu1987):

Move all terms to left side and take 3 as common factors from all the terms...

OpenStudy (anonymous):

\(3x^2=-21x-15\) 1st of all bring 21x to the left of equation

OpenStudy (anonymous):

\(3x^2-21x=-15\) now bring -15 to the left side and it would change into postive as \(3x^2-21x+15\)

OpenStudy (anonymous):

now factorize them @chariety101 understanding?

OpenStudy (anonymous):

multiply 15 to 3 the first and the end co-efficients with each other

OpenStudy (anonymous):

15*3=45 now take LCM of 45 as 

OpenStudy (anonymous):

alright alright i think i understand... so far..

OpenStudy (anonymous):

are you there? @ziqbal103

OpenStudy (anonymous):

@chariety101 i think in your question 3*2 is ...isn't it?

OpenStudy (anonymous):

i did wrong by assuming that \(3x^2\) but it is 3*2=6 @chariety101

OpenStudy (midhun.madhu1987):

 now solve this quadratic equation by quadratic formula... do u know the quadratic formula?